PROGRESS ON THE INFRARED SPECTROSCOPY OF DIAMONDOIDS AND THEIR ASTROPHYSICAL APPLICATION

\underline{O. PIRALI}, J. OOMENS, N. POLFER, FOM Rijnhuizen, 3439MN Nieuwegein, The Netherlands; Y. UENO, R. MABOUDIAN, Department of Chemical Engineering, U.C. Berkeley, USA; P. MAY, J. FILIK, School of Chemistry, University of Bristol, UK; J. DAHL, S. LIU, B. CARLSON, Molecular Diamond Technologies, ChevronTexaco Technology Ventures, Richmond, USA.

In the continuous effort for the unambiguous assignment of large carbonaceous species, we will present our recent laboratory observations concerning the infrared spectra of higher diamondoids (molecules consisting of diamond-like carbon cages) obtained by using various spectroscopic techniques. Increasing the size of the sample from adamantane (C_{10}H_{16}) to diamantane (C_{14}H_{20}), triamantane (C_{18}H_{24}), etc, corresponds to an addition of a cage C_{4}H_{4} entity. Beyond its very interesting spectroscopic interest (due to the high symmetry of the systems studied here), this family of compounds which has been found in meteorites^{a}, also display a strong vibrational transition around 3.5 microns, thereby providing a convincing agreement with observed interstellar emission bands^{b}. We will discuss in this communication the spectroscopic analysis of our different spectra as well as our search for the possible occurrence of these compounds in several astrophysical observations.

aE. Anders and E. Zinner, Meteorites 28 (1993) 490

bL. J. Allamandola, S.A. Sandford, A.G .G. M. Tielens, T. M. Herbst, Astrophysical Journal 399 (1992) 134